Consider the lines through P(2,4) and perpendicular to the x- and y-axes, respectively. Find the area and the perimeter of the rectangle formed by these lines and the axes.
I'm totally lost, please help. Thanks so much.

Answer by richwmiller(17219)

Ok I think I understand what it wants.
We are using the x axis and the y axis as two sides boundaries.
We are using the line segment parallel to the x axis that passes through 2,4 as another side
and the line segment that is parallel to the y axis that passes though 2,4 as the other side.

so the points of the corners are (2,4) (0,4) (0,0) and (2,0)
So the lengths of the sides are 2 and 4
perimeter = 2*(2+4)=12
area = 2*4=8